An IR system returns 30 relevant documents, and 17 irrelevant documents. There is a total of 90 relevant documents in the collection. What is the precision of the system on this search, and what is its recall (include the calculation steps in your answer)?
Answer #1

Relevant documents = 30

Irrelevant documents = 17

Total relevant documents in collection = 90

Therefore, Precision = 30/(30+17) = 30/47 = 0.638

and Recall = 30 / 90 = 0.333